OpenStudy (hanna1):

PLZ HELP WILL MEDAL AND FAN!! Find the slope of the line whose equation is 2x -4y =10

OpenStudy (hanna1):

-1/2 ½ 4 -2

OpenStudy (will.h):

2x -4y = 10 Set that in slope intercept form Y= mx+b Where m is the slope -4y = -2x + 10 Y = (2/4)x -10/4 Y = (1/2)x - 10/4 Therefore slope is 1/2